OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -5) with a slope of -3. y = -3x - 13 y = -3x + 11 y = -3x + 13 y = -3x + 1

OpenStudy (anonymous):

@Jhannybean

OpenStudy (bahrom7893):

y - y1 = M(x - x1) y - (-5) = -3(x - 2) Simplify

OpenStudy (anonymous):

y + 5 = -6x + 12

OpenStudy (anonymous):

y = -6x -7 ?

OpenStudy (eli_moses):

????? umm...... no they give you already the slope, right?

OpenStudy (anonymous):

yes

OpenStudy (eli_moses):

so you already have y=-3x+b

OpenStudy (anonymous):

yes

OpenStudy (eli_moses):

they give you the point (2, -5) so you substitute it into the equation. Do you know how to do that?

OpenStudy (anonymous):

ik of it but im not sure how to do it

OpenStudy (eli_moses):

ok i'll show you

OpenStudy (anonymous):

k

OpenStudy (eli_moses):

so you have \[y=-3x+b\] now you plug in the coordinates (2, -5). The equation changes to: \[-5=-3(2)+b\] Now you're finding what b is. Does this help a bit?

OpenStudy (anonymous):

Yes, Do i subtract -5?

OpenStudy (eli_moses):

No. You're trying to make each side balanced. First multiply -3 and 2. Try making both sides balanced.

OpenStudy (anonymous):

-5 = -6x + b

OpenStudy (eli_moses):

Yes. Now divide both sides by -6.

OpenStudy (eli_moses):

To find b you have to single it out with a number on the opposite side of the equation.

OpenStudy (anonymous):

-5/6 = x + b

OpenStudy (eli_moses):

You don't need the x since you substituted it.

OpenStudy (anonymous):

what is equation looking like right now

OpenStudy (anonymous):

you make it look like -5/6 = b

OpenStudy (eli_moses):

it's -5/-6=b

OpenStudy (eli_moses):

wait, there might be something wrong......

OpenStudy (eli_moses):

oh..... so we're back to the equation -5=-6+b

OpenStudy (eli_moses):

you add 6 to both sides since it's the inverse. I got confused because you put the x in front of the -6 before.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y = -3x + 1

OpenStudy (anonymous):

D?

OpenStudy (eli_moses):

Yes. Exactly. To check your answer just plug in the coordinates into the equation you got and see if it equals itself.

OpenStudy (anonymous):

ok thanks

OpenStudy (eli_moses):

no problem